I need a spring to move an object 10 inches in the air when activated. The spring needs to be 2 to 3 inches in diameter. The object will weigh 6 ounces. The object will roll in to a hole and contact the spring. How can I buy a spring with these specs? Thanks!
1
Expert's answer
2012-10-01T10:13:18-0400
ma=k|x1+x|-mg
ma=k|mg/k+x|-mg=k*|x|
k=1
x=10+(3-2)=11
m=6
ma=kx
a=11/6